Dublin University Rifle Club

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Mark Dennehy 167 Applewood Hts Greystones Co. Wicklow Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Dave Cooney Ardnaglass Corrig Road Dun Laoghaire

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

John K. Walsh Carjon Lodge Cross Chapel Blessington Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Agnes Graham 114 Towerview Crescent Bangor Co. Down BT19 6A2

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Tom McKnight 26 Meadow Drive Ballyclare County Antrim BT39 9XF Northern Ireland

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Karen Farrell 41 Deerpark Walk Blessington Co. Wicklow

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,

DURC · c/o DUCAC · House 27 · Trinity College Dublin · Dublin 2 · Ireland

Geoff Cooney Wilkinstown Navan Co. Meath

Mark Dennehy, DURC Secretary Vision Lab O’Reilly Institute Trinity College Dublin Dublin 2

☎: +353-1-6081281

Fax: +353-1-6772204

e-mail:

Mark.Dennehy@cs.tcd.ie

Subject: Trinity Top Ten  27th July 2001

A Chara,

Well, we’ve had some problems in the starting of the TTT this year, with the delayed DURC Open stomping all over the starting date for the TTT and delaying me in printing stickers and distributing them. In order to compensate, the deadlines were pushed back and those now printed on the enclosed stickers are accurate. However, since the Round One deadline is upon us, it will not be enforced, so the round one cards are due at the same time as the round two cards. That is, August 10 is the deadline for rounds one and two.

Further, you will note that there are more stickers per round for different leagues - there are 20 shots per round, regardless of league, so for 50m cards there is only one sticker required. For 25yd target cards (10-bull), there are two stickers per round as you must shoot two cards, and for Air Rifle’s 5-bull targets, there are four stickers per round.

If there are any queries, please don’t hesitate to ask me at the above contact numbers or email.

Good luck to all.

Misé le meas,